# Sample-Efficient Reinforcement Learning from Human Feedback via Information-Directed Sampling

### v1

@@ IV-B3 Construct the surrogate environment
IV-B4 IDS with surrogate environments
The pseudo-code of our IDS algorithm (with the learning target being the surrogate environment ~$\tilde{\mathcal{E}}^{*}_{t}$) is shown in Algorithm 1. Roughly speaking, the agent, at each episode $t$, first computes the posterior distributions of the transition kernel $P$ and reward function $R$ (as shown in Eqns. (14)-(15)). Then, the agent computes the surrogate environment ~$\tilde{\mathcal{E}}^{*}_{t}$ based on Eqn. (10), and chooses the policy
$\pi_{\mathrm{IDS}}^{t}=\arg\max_{\pi\in\Pi}\mathbb{E}_{t}[V_{1,\pi}^{\mathcal{E}}(s_{1})]+\frac{\lambda}{2}\cdot\mathbb{I}_{t}^{\pi}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right).$
@@ IV-B4 IDS with surrogate environments
IV-C Regret analysis of Algorithm 1
@@
Since the reward is bounded by $[0,1]$, we have $\alpha_{\mathcal{E}}\leq H+1$. The average value diameter over $\Theta$ is denoted by $\alpha\triangleq\mathbb{E}_{\mathcal{E}\sim\rho}\big{[}\alpha_{\mathcal{E}}^{2}\big{]}^{1/2}$. Similar to the prior work [7], we need to make the following assumption about posterior consistency.
Assumption 3 (Posterior Consistency).
@@
This means as the sample size approaches infinity, the posterior distribution of environment obtained through Eqns. (14)-(15) tends to concentrate around the true distribution. In other words, the posterior distribution will correctly identify the true environment that generates these trajectory data.
Theorem 1.
@@
$BR_{T}(\pi_{\text{IDS}})\leq\alpha\sqrt{TH\log(K(\epsilon))}+T\epsilon+T_{0},$ (16)
@@
The detailed proof of Theorem 1 is deferred to Appendix A-A. We point out that the existing TS-based RLHF algorithm [8] has an upper bound of order
$\tilde{O}(H^{2}\sqrt{T(\ell_{P}+\ell_{R})}(\text{dim}_{1}(P,1/T))+\text{dim}_{1}(R,1/T)),$
@@
Remark 3.
The regret upper bounds in some related works [35, 8] are related to the derivative bound of the link function. However, our upper bound is independent of the link function we use (which is the sigmoid function). This is because the posterior consistency assumption implicitly imposes requirements on the link function — the link function should be monotonically increasing to ensure that better trajectories correspond to higher preference probabilities. For example, if the link function is equal to a constant $\frac{1}{2}$, then the posterior distribution of rewards would not change (according to the posterior update rule in Eqn. (15)), and thus could not converge to the true distribution. Therefore, the assumption in previous work of a strictly positive lower bound on the link function’s derivative is encompassed by our posterior consistency assumption.
IV-D Applications
@@
Definition 3 (Tabular RLHF).
@@
Definition 4 (Linear RLHF).
Let $\phi^{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ and $\phi^{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ be known feature maps with bounded norms $||\phi^{P}(s,a)||_{2}\leq 1$ and $||\phi^{R}(s,a)||_{2}\leq 1$. We say a Bayesian RLHF problem is linear if for any =$\mathcal{E}=\{(P_{h}^{\mathcal{E}},R_{h}^{\mathcal{E}})\}_{h=1}^{H}\in\Theta$, there exists vector-valued maps $\psi_{h}^{P,\mathcal{E}}$ and $\psi_{h}^{R,\mathcal{E}}$ with bounded $\ell_{2}$-norm such that for every $(s,a)\in\mathcal{S}\times\mathcal{A}$,
@@
Specializing Theorem 1 to tabular and linear Bayesian RLHF problems, we have the following Bayesian regret bounds. The proofs of Theorems 2 and 3 are deferred to Appendices A-B and A-C respectively.
Theorem 2 (Tabular RLHF).
@@
where $T_{0}$ is a fixed integer that is independent of $T$. Setting $\epsilon=\frac{1}{T}$, our regret bound is of order $\tilde{O}(\sqrt{SAH^{4}T})$.
Recall that the IDS algorithm proposed for the tabular RL setting [9] has a regret upper bound of order $\tilde{O}(\sqrt{S^{2}A^{2}H^{4}T})$. Compared to their result, our method relies on less informative data (preference feedback instead of directly observable rewards) but achieves a better regret bound by a factor of $S$ and $A$. This improvement is primarily due to our refined analytical techniques, inspired by recent advancements in TS [7].
Theorem 3 (Linear RLHF).
@@
$BR_{T}(\pi_{\mathrm{IDS}})\leq\alpha H\sqrt{dT\log(K_{\mathcal{F}}(\epsilon))}+T\epsilon+T_{0},$ (17)
@@
Compared to [8], which derives a regret upper bound of $\tilde{O}(H^{11/2}d^{17/2}\sqrt{T})$ for their TS algorithm, our regret upper bound is better when the covering number of the linear MDP is not of exponential size. If we convert their result to the tabular setting by coarsely substituting $d$ with $SA$, our regret bound is also better in terms of $H,S,A.$ However, we also point out that the above comparison is not an apples-to-apples comparison, as we consider Bayesian regret, while they consider frequentist regret, and their algorithm also accounts for the number of queries.
Corollary 1 (Contextual Dueling Bandits).
Contextual dueling bandits are a simplified version of our MDP setting (with $H=1$) and have been extensively studied in previous RLHF research [6, 2, 23]. By setting $H=1$ in Theorem 3, the Bayesian regret for Algorithm 1 in the linear contextual dueling bandit problem satisfies
@@
for any $\epsilon>0$ and sufficiently large $T$. Setting $\epsilon=\frac{1}{T}$, the regret upper bound is of order $\tilde{O}(\sqrt{dT})$.
Without considering the covering number of linear environment, our regret upper bound is better than $\tilde{O}(d\sqrt{T})$ derived by [23]. Another work [20] assumes a finite number of arms with a regret upper bound of $\tilde{O}(\sqrt{dT})$, while we assume that the parameter space of the linear MDP is compact.
V The Approximate-IDS Algorithm
While the IDS algorithm (Algorithm 1) is principled and sample-efficient, it suffers from relatively high computational complexity. This is because the calculation of the surrogate environment ~$\tilde{\mathcal{E}}_{t}^{*}$ (which depends on the construction of $\epsilon$-value partition) is challenging. As a remedy, we develop a computationally efficient algorithm, named Approximate-IDS, whose optimization objective is independent of the surrogate environment (thus avoids the partition of $\Theta$ in computation) and has finer properties for analysis. This allows the algorithm to be computed efficiently by traditional RL algorithms from standard RL theory.
V-1 Algorithm description
@@
$\mathrm{KL}^{h}_{s,a}(\mathcal{E},\mathcal{E}^{\prime})\!\triangleq\!D_{\mathrm{KL}}\big{(}(P_{h}^{\mathcal{E}}\otimes R_{h}^{\mathcal{E}})(\cdot|s,a)||(P_{h}^{\mathcal{E}^{\prime}}\otimes R_{h}^{\mathcal{E}^{\prime}})(\cdot|s,a)\big{)},$
$\bar{r}_{h}(s_{h},a_{h})\triangleq r_{h}(s_{h},a_{h})+\frac{\lambda}{2}\cdot\mathbb{E}_{t}\big{[}\mathrm{KL}_{s_{h},a_{h}}^{h}(\mathcal{E},\bar{\mathcal{E}}_{t})\big{]},$
where ¯$\bar{\mathcal{E}}_{t}$ denotes the posterior mean of $\mathcal{E}$ given $\mathcal{D}_{t}$, i.e., $P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|s,a)=\mathbb{E}_{t}[P_{h}^{\mathcal{E}}(\cdot|s,a)]$ and $R_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|s,a)=\mathbb{E}_{t}[R_{h}^{\mathcal{E}}(\cdot|s,a)]$.
The overall procedure is similar to that of Algorithm 1, with the key difference being the selection of the IDS policy (Line 4). Intuitively, $\mathcal{E}$ is sufficiently close to ~$\tilde{\mathcal{E}}_{t}^{*}$ under metric $\ell_{g}$, thus it is reasonable to use $\mathcal{E}$ directly for mutual information computation. Given trajectories and rewards, the additional environmental information revealed by human feedback, i.e., $\mathbb{I}_{t}^{\pi}\big{(}\tilde{\mathcal{E}}_{t}^{*};o_{t}\mid(\mathcal{H}_{t,H},\mathcal{R}_{t,H})\big{)}$, can be disregarded. Thus, we use the entire environment $\mathcal{E}$ instead of the surrogate environment ~$\tilde{\mathcal{E}}_{t}^{*}$ to compute the mutual information and discard the information of the trajectory generated by the baseline policy $\pi_{0}$ and human feedback. Therefore, we replace the mutual information term $\mathbb{I}_{t}^{\pi}\big{(}\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\big{)}$ by $\sum_{h=1}^{H}\mathbb{E}_{t}\big{[}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}[\mathrm{KL}_{s_{h},a_{h}}^{h}(\mathcal{E},\bar{\mathcal{E}}_{t})]\big{]}$ (Eqn. (62) in Lemma 5). We can compute the approximate IDS policy as follows:
$\pi_{\mathrm{app}}^{t}$ $\!=\!\arg\max_{\pi\in\Pi}\mathbb{E}_{t}[V_{1,\pi}^{\mathcal{E}}(s_{1})]\!+\!\frac{\lambda}{2}\sum_{h=1}^{H}\mathbb{E}_{t}\big{[}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}[\mathrm{KL}_{s_{h},a_{h}}^{h}(\mathcal{E},\bar{\mathcal{E}}_{t})]\big{]}$ (18)
$=\arg\max_{\pi\in\Pi}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\left[\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\right].$
@@
Note that $\bar{r}$ and ¯$\bar{\mathcal{E}}_{t}$ are both independent of the surrogate environment, and can be well approximated by Monte Carlo sampling. Therefore, by introducing $\bar{r}$, solving $\pi_{\mathrm{app}}^{t}$ at episode $t$ is equivalent to finding an optimal policy based on MDP $\{P_{h}^{\bar{\mathcal{E}}_{t}},\bar{r}_{h}\}_{h=1}^{H}$, which can be solved efficiently by the PPO algorithm [14].
V-2 Regret bounds for Approximate-IDS
@@
Proposition 1.
We define
$r^{\prime}_{h}(s,a)\triangleq r_{h}(s,a)+\frac{\lambda}{2}\mathbb{E}_{t}\big{[}\mathrm{KL}_{s,a}^{h}(\tilde{\mathcal{E}}_{t}^{*},\bar{\mathcal{E}}_{t})\big{]}.$ (19)
@@
$\bigg{|}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r^{\prime}_{h}(s_{h},a_{h})\bigg{]}-\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}\bigg{|}\leq\frac{\lambda}{2}\epsilon(1-2\log\beta).$ (20)
The proof is deferred to Appendix A-D. To better understand this proposition, consider an extreme scenario: we divide the environment into the smallest units, with each $\Theta_{k}^{\epsilon}$ containing only one environment. We have =~$\mathcal{E}=\tilde{\mathcal{E}}_{t}^{*}$. The left hand of Eqn. (20) equals 00, so Proposition 1 holds true. Since $|a-b|\leq B|\log\frac{a}{b}|$ for any $a,b\in(0,B)$, we have $\ell_{1}(P,Q)\leq B\ell_{g}(P,Q)$. If we ignore the constant $B$, by fixing the $\epsilon$-value, our distance achieves a finer environmental partition. On this finer partition, $\mathcal{E}$ and ~$\tilde{\mathcal{E}}_{t}^{*}$ behave more similarly, allowing us to ensure that Proposition 1 holds. Then, using Proposition 1, we give the Bayesian regret bound for the Approximate-IDS algorithm.
Theorem 4.
@@
$BR_{T}(\pi_{\mathrm{app}})\leq\alpha\sqrt{2TH\log(K(\epsilon))}+\left(1+\frac{(1-2\log\beta)}{2}\sqrt{\frac{\alpha^{2}TH}{2\log(K(\epsilon))}}\right)T\epsilon+T_{0}.$ (21)
By choosing a small $\epsilon$, the regret upper bound is of order $O\big{(}\sqrt{H^{3}T\log(K(\epsilon))}\big{)}$, matching that of Algorithm 1 presented in Sec. IV-C.
@@
By the optimality of $\pi_{\mathrm{app}}$, we have
$\mathbb{E}_{\pi_{\mathrm{app}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}$ $\geq\mathbb{E}_{\pi_{\mathrm{IDS}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}$
$\geq\mathbb{E}_{\pi_{\mathrm{IDS}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r_{h}(s_{h},a_{h})\bigg{]}$
$=\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{IDS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right].$ (22)

### v3

@@ 4.2.3 Construct the surrogate environment
4.2.4 IDS with surrogate environments
The pseudo-code of our IDS algorithm (with the learning target being the surrogate environment $\widetilde{\mathcal{E}}^{*}_{t}$) is shown in Algorithm 1. The algorithm requires priors $\rho^{P},\rho^{R}$ as input. Prior refers to the initial assumptions about model parameters before learning begins. Suitable priors can be derived from existing domain knowledge (e.g., robot dynamics parameters, user behavior patterns). For instance, a Gaussian prior might be adopted in robotic  Haninger et al. (2022, 2023), while a Dirichlet prior could be used in multi-agent collaboration  Wu et al. (2021). In our paper, an appropriate prior can be selected for the IDS algorithm according to practical applications. Roughly speaking, the agent, at each episode $t$, first computes the posterior distributions of the transition kernel $P$ and reward function $R$ (as shown in Eqns. (4.13)-(4.14)). Then, the agent computes the surrogate environment $\widetilde{\mathcal{E}}^{*}_{t}$ based on Eqn. (4.9), and chooses the policy
$\pi_{\mathrm{IDS}}^{t}=\arg\max_{\pi\in\Pi}\mathbb{E}_{t}[V_{1,\pi}^{\mathcal{E}}(s_{1})]+\frac{\lambda}{2}\cdot\mathbb{I}_{t}^{\pi}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right).$
@@ 4.2.4 IDS with surrogate environments
4.3 Regret analysis of Algorithm 1
@@
Since the reward is bounded by $[0,1]$, we have $\alpha_{\mathcal{E}}\leq H+1$. The average value diameter over $\Theta$ is denoted by $\alpha\triangleq\mathbb{E}_{\mathcal{E}\sim\rho}\big{[}\alpha_{\mathcal{E}}^{2}\big{]}^{1/2}$. Similar to the prior work Moradipari et al. (2023), we need to make the following assumption about posterior consistency.
Assumption 4.10 (Posterior Consistency).
@@
This means as the sample size approaches infinity, the posterior distribution of environment obtained through Eqns. (4.13)-(4.14) tends to concentrate around the true distribution. In other words, the posterior distribution will correctly identify the true environment that generates these trajectory data.
Theorem 4.11.
@@
$BR_{T}(\pi_{\text{IDS}})\leq\alpha\sqrt{TH\log(K(\epsilon))}+T\epsilon+T_{0},$ (4.15)
@@
The detailed proof of Theorem 4.11 is deferred to Appendix A.1. We point out that the existing TS-based RLHF algorithm Wu and Sun (2023) has an upper bound of order
$\widetilde{O}(H^{2}\sqrt{T(\ell_{P}+\ell_{R})}(\text{dim}_{1}(P,1/T))+\text{dim}_{1}(R,1/T)),$
@@
Remark 4.12.
The regret upper bounds in some related works Saha et al. (2023); Wu and Sun (2023) are related to the derivative bound of the link function. However, our upper bound is independent of the link function we use (which is the sigmoid function). This is because the posterior consistency assumption implicitly imposes requirements on the link function — the link function should be monotonically increasing to ensure that better trajectories correspond to higher preference probabilities. For example, if the link function is equal to a constant $\frac{1}{2}$, then the posterior distribution of rewards would not change (according to the posterior update rule in Eqn. (4.14)), and thus could not converge to the true distribution. Therefore, the assumption in previous work of a strictly positive lower bound on the link function’s derivative is encompassed by our posterior consistency assumption.
4.4 Applications
@@
Definition 4.13 (Tabular RLHF).
@@
Definition 4.14 (Linear RLHF).
Let $\phi^{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ and $\phi^{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ be known feature maps with bounded norms $\phi^{P}(s,a)||_{2}\leq 1$ and $\phi^{R}(s,a)||_{2}\leq 1$. We say a Bayesian RLHF problem is linear if for any $\mathcal{E}=\{(P_{h}^{\mathcal{E}},R_{h}^{\mathcal{E}})\}_{h=1}^{H}\in\Theta$, there exists vector-valued maps $\psi_{h}^{P,\mathcal{E}}$ and $\psi_{h}^{R,\mathcal{E}}$ with bounded $\ell_{2}$-norm such that for every $(s,a)\in\mathcal{S}\times\mathcal{A}$,
@@
Specializing Theorem 4.11 to tabular and linear Bayesian RLHF problems, we have the following Bayesian regret bounds. The proofs of Theorems 4.15 and 4.16 are deferred to Appendices A.2 and A.3 respectively.
Theorem 4.15 (Tabular RLHF).
@@
where $T_{0}$ is a fixed integer that is independent of $T$. Setting $\epsilon=\frac{1}{T}$, our regret bound is of order $\widetilde{O}(\sqrt{SAH^{4}T})$.
Recall that the IDS algorithm proposed for the tabular RL setting Hao and Lattimore (2022) has a regret upper bound of order $\widetilde{O}(\sqrt{S^{2}A^{2}H^{4}T})$. Compared to their result, our method relies on less informative data (preference feedback instead of directly observable rewards) but achieves a better regret bound by a factor of $S$ and $A$. This improvement is primarily due to our refined analytical techniques, inspired by recent advancements in TS Moradipari et al. (2023).
Theorem 4.16 (Linear RLHF).
@@
$BR_{T}(\pi_{\mathrm{IDS}})\leq\alpha H\sqrt{dT\log(K_{\mathcal{F}}(\epsilon))}+T\epsilon+T_{0},$ (4.16)
@@
Compared to Wu and Sun (2023), which derives a regret upper bound of $\widetilde{O}(H^{11/2}d^{17/2}\sqrt{T})$ for their TS algorithm, our regret upper bound is better when the covering number of the linear MDP is not of exponential size. If we convert their result to the tabular setting by coarsely substituting $d$ with $SA$, our regret bound is also better in terms of $H,S,A.$ However, we also point out that the above comparison is not an apples-to-apples comparison, as we consider Bayesian regret, while they consider frequentist regret, and their algorithm also accounts for the number of queries.
Corollary 4.17 (Contextual Dueling Bandits).
Contextual dueling bandits are a simplified version of our MDP setting (with $H=1$) and have been extensively studied in previous RLHF research Ye et al. (2024); Zhu et al. (2023); Li et al. (2024). By setting $H=1$ in Theorem 4.16, the Bayesian regret for Algorithm 1 in the linear contextual dueling bandit problem satisfies
@@
for any $\epsilon>0$ and sufficiently large $T$. Setting $\epsilon=\frac{1}{T}$, the regret upper bound is of order $\widetilde{O}(\sqrt{dT})$.
Without considering the covering number of linear environment, our regret upper bound is better than $\widetilde{O}(d\sqrt{T})$ derived by Li et al. (2024). Another work Saha (2021) assumes a finite number of arms with a regret upper bound of $\widetilde{O}(\sqrt{dT})$, while we assume that the parameter space of the linear MDP is compact.
5 The Approximate-IDS Algorithm
While the IDS algorithm (Algorithm 1) is principled and sample-efficient, it suffers from relatively high computational complexity. This is because the calculation of the surrogate environment $\widetilde{\mathcal{E}}_{t}^{*}$ (which depends on the construction of $\epsilon$-value partition) is challenging. As a remedy, we develop a computationally efficient algorithm, named Approximate-IDS, whose optimization objective is independent of the surrogate environment (thus avoids the partition of $\Theta$ in computation) and has finer properties for analysis. This allows the algorithm to be computed efficiently by traditional RL algorithms from standard RL theory.
5.1 Algorithm description
@@
$\mathrm{KL}^{h}_{s,a}(\mathcal{E},\mathcal{E}^{\prime})\!\triangleq\!D_{\mathrm{KL}}\big{(}(P_{h}^{\mathcal{E}}\otimes R_{h}^{\mathcal{E}})(\cdot|s,a)||(P_{h}^{\mathcal{E}^{\prime}}\otimes R_{h}^{\mathcal{E}^{\prime}})(\cdot|s,a)\big{)},$ (5.1)
and
$\bar{r}_{h}(s_{h},a_{h})\triangleq r_{h}(s_{h},a_{h})+\frac{\lambda}{2}\cdot\mathbb{E}_{t}\big{[}\mathrm{KL}_{s_{h},a_{h}}^{h}(\mathcal{E},\bar{\mathcal{E}}_{t})\big{]},$ (5.2)
where $\bar{\mathcal{E}}_{t}$ denotes the posterior mean of $\mathcal{E}$ given $\mathcal{D}_{t}$, i.e., $P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|s,a)=\mathbb{E}_{t}[P_{h}^{\mathcal{E}}(\cdot|s,a)]$ and $R_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|s,a)=\mathbb{E}_{t}[R_{h}^{\mathcal{E}}(\cdot|s,a)]$.
The overall procedure is similar to that of Algorithm 1, with the key difference being the selection of the IDS policy (Line 4). Intuitively, $\mathcal{E}$ is sufficiently close to $\widetilde{\mathcal{E}}_{t}^{*}$ under metric $\ell_{g}$, thus it is reasonable to use $\mathcal{E}$ directly for mutual information computation. Given trajectories and rewards, the additional environmental information revealed by human feedback, i.e., $\mathbb{I}_{t}^{\pi}\big{(}\widetilde{\mathcal{E}}_{t}^{*};o_{t}\mid(\mathcal{H}_{t,H},\mathcal{R}_{t,H})\big{)}$, can be disregarded. Thus, we use the entire environment $\mathcal{E}$ instead of the surrogate environment $\widetilde{\mathcal{E}}_{t}^{*}$ to compute the mutual information and discard the information of the trajectory generated by the baseline policy $\pi_{0}$ and human feedback. Therefore, we replace the mutual information term $\mathbb{I}_{t}^{\pi}\big{(}\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\big{)}$ by $\sum_{h=1}^{H}\mathbb{E}_{t}\big{[}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}[\mathrm{KL}_{s_{h},a_{h}}^{h}(\mathcal{E},\bar{\mathcal{E}}_{t})]\big{]}$ (Eqn. (C.4) in Lemma C.2). We can compute the approximate IDS policy as follows:
$\pi_{\mathrm{app}}^{t}$ $\!=\!\arg\max_{\pi\in\Pi}\mathbb{E}_{t}[V_{1,\pi}^{\mathcal{E}}(s_{1})]\!+\!\frac{\lambda}{2}\sum_{h=1}^{H}\mathbb{E}_{t}\big{[}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}[\mathrm{KL}_{s_{h},a_{h}}^{h}(\mathcal{E},\bar{\mathcal{E}}_{t})]\big{]}$ (5.3)
$\overset{(a)}{=}\arg\max_{\pi\in\Pi}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\left[\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\right],$
where (a) uses the linearity of expectation and independence of priors over different layers.
@@
Note that $\bar{r}$ and $\bar{\mathcal{E}}_{t}$ are both independent of the surrogate environment, and can be well approximated by Monte Carlo sampling. Therefore, by introducing $\bar{r}$, solving $\pi_{\mathrm{app}}^{t}$ at episode $t$ is equivalent to finding an optimal policy based on MDP $\{P_{h}^{\bar{\mathcal{E}}_{t}},\bar{r}_{h}\}_{h=1}^{H}$, which can be solved efficiently by the PPO algorithm Schulman et al. (2017).
5.2 Regret bounds for Approximate-IDS
@@
Proposition 5.1.
For any $\epsilon$-value partition $\{\Theta_{k}^{\epsilon}\}_{k=1}^{K(\epsilon)}$ (Eqn. (4.7)) and the surrogate environment $\widetilde{\mathcal{E}}_{t}^{*}$ constructed by Eqn. (4.9), Define
$r^{\prime}_{h}(s,a)\triangleq r_{h}(s,a)+\frac{\lambda}{2}\mathbb{E}_{t}\big{[}\mathrm{KL}_{s,a}^{h}(\widetilde{\mathcal{E}}_{t}^{*},\bar{\mathcal{E}}_{t})\big{]}.$ (5.4)
@@
$\bigg{|}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r^{\prime}_{h}(s_{h},a_{h})\bigg{]}-\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}\bigg{|}\leq\frac{\lambda}{2}\epsilon(1+\log\frac{B}{\beta}).$ (5.5)
The proof is deferred to Appendix A.4. The proof indicates that the proposition may fail to hold when the surrogate environment is constructed using the $\ell_{1}$-distance or KL divergence. To better understand this proposition, consider an extreme scenario: we divide the environment into the smallest units, with each $\Theta_{k}^{\epsilon}$ containing only one environment. We have $\mathcal{E}=\widetilde{\mathcal{E}}_{t}^{*}$. The left hand of Eqn. (5.5) equals 0, so Proposition 5.1 holds true. Since $|a-b|\leq B|\log\frac{a}{b}|$ for any $a,b\in(0,B)$, we have $\ell_{1}(P,Q)\leq B\ell_{g}(P,Q)$. If we ignore the constant $B$, by fixing the $\epsilon$-value, our distance achieves a finer environmental partition. On this finer partition, $\mathcal{E}$ and $\widetilde{\mathcal{E}}_{t}^{*}$ behave more similarly, allowing us to ensure that Proposition 5.1 holds. Then, using Proposition 5.1, we give the Bayesian regret bound for the Approximate-IDS algorithm.
Theorem 5.2.
@@
$BR_{T}(\pi_{\mathrm{app}})\leq\alpha\sqrt{2TH\log(K(\epsilon))}+\left(1+\frac{(1+\log(B/\beta))}{2}\sqrt{\frac{\alpha^{2}TH}{2\log(K(\epsilon))}}\right)T\epsilon+T_{0}.$ (5.6)
By choosing a small $\epsilon$, the regret upper bound is of order $O\big{(}\sqrt{H^{3}T\log(K(\epsilon))}\big{)}$, matching that of Algorithm 1 presented in Sec. 4.3.
@@
First notice that
$\mathbb{E}_{\pi_{\mathrm{app}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}\overset{(a)}{\geq}\mathbb{E}_{\pi_{\mathrm{IDS}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}\overset{(b)}{\geq}\mathbb{E}_{\pi_{\mathrm{IDS}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r_{h}(s_{h},a_{h})\bigg{]}\overset{(c)}{=}\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{IDS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right],$
where (a) uses the optimality of $\pi_{\mathrm{app}}^{t}$, (b) follows from Eqn. (5.2),(c) uses the definition of $\bar{\mathcal{E}}_{t}$ and the linearity of expectation.